### v3

# Probing Einstein-Maxwell-Scalar Black hole via Thin Accretion Disks and Shadows with EHT Observations of M87* and Sgr A*

Abstract
[…]
I Introduction
Black holes (BHs) represent one of the fundamental solutions to Einstein’s field equations and are regarded as robust predictions of General Relativity (GR). While historically debated, numerous observational efforts have been made to confirm their existence. A significant breakthrough occurred in 2015 when the Laser Interferometer Gravitational-Wave Observatory (LIGO) detected gravitational waves (GWs) originating from a binary BH merger 124 . Subsequently, in 2019, the Event Horizon Telescope (EHT) provided the first direct observational evidence of a BH shadow at the center of the galaxy M87 125 ; 126 ; 127 ; 128 ; 129 ; 130 .
Further advancements in EHT observations led to the polarization imaging of M87*, revealing the presence of a magnetic field, which provides insights into jet formation mechanisms 131 ; 132 ; 133 . Additionally, EHT reported the detection of the shadow of SgrA*, the supermassive BH at the center of the Milky Way 134 . These observations offer valuable constraints on the geometry of spacetime in the strong-field regime, particularly near the event horizon, allowing precise measurements of BH mass and spin 135 ; 136 ; 145 ; 146 ; 147 ; 148 ; 149 ; 150 ; 151 ; 152 ; 153 ; 154 ; 155 ; 185 . Collectively, these findings provide a crucial platform for testing GR and exploring potential modifications to gravitational theories in extreme astrophysical environments 137 ; Liu:2024brf ; Yin:2025coq ; He:2022lrc ; Wu:2024sng ; Feng:2024iqj ; Feng:2023iha ; Feng:2022bst .
It is well known that GR has demonstrated remarkable accuracy in various tests 138 . However, it faces certain limitations within the cosmological framework. These include challenges in explaining the accelerated expansion of the universe, the presence of initial singularities, the missing mass problem, and the Coincidence Problem. Furthermore, testing GR in the strong-field regime, particularly in the vicinity of BH horizons, remains a significant challenge 139 . In addition, modifications to GR have been explored to extend our understanding beyond its conventional framework, aiming to probe near-horizon physics and the large-scale structure of the universe 140 ; 141 ; 142 ; 180 .
Among various modified gravity models, the Einstein-Maxwell-Scalar (EMS) model serves as a promising candidate. In fact, various specific realizations of EMS models naturally arise as consistent truncations or low-energy limits of string theory. For instance, in heterotic string theory, compactification from ten to four dimensions in the Einstein frame yields an effective action containing the metric, a dilaton-like scalar field, and Abelian gauge fields with an exponential coupling of the form $f(\phi)F_{\mu\nu}F^{\mu\nu}$ — as in the well-known Gibbons–Maeda–Garfinkle–Horowitz–Strominger (GMGHS) solution Gibbons:1988 ; Garfinkle:1990qj . Similar structures occur in Type II string compactifications, where moduli fields (Kähler, complex structure, axion–dilaton) couple to gauge fields descending from the NS–NS and R–R sectors Polchinski:1998 ; Becker:2006 . In extended supergravity, particularly $\mathcal{N}=4$ and $\mathcal{N}=8$ theories, the gauge kinetic matrix $f_{IJ}(\phi)$ depends explicitly on scalar fields Cremmer:1979 ; deWit:1982 . Truncation to a single $U(1)$ vector and one scalar degree of freedom yields precisely the EMS form. The $\mathcal{N}=2$ STU model Duff:1996 provides another example, in which selecting a single scalar–vector sector produces an EMS-type action.
The EMS action serves as a powerful unifying framework, encompassing two distinct and mutually exclusive branches of black hole solutions: the Einstein–Maxwell–Dilaton (EMD) branch and the spontaneously scalarized Reissner–Nordström (RN) branch. The profound differences between these two scenarios are encoded solely in the local behavior of the coupling function $K(\phi)$ at $\phi=0$. The EMD limit, realized by choosing an exponential coupling $K(\phi)=\mathrm{e}^{2\alpha\phi}$, is characterized by $K(0)=1$ and $K^{\prime}(0)\neq 0$. This condition renders the electro-vacuum RN solution, where $\phi\equiv 0$, an invalid background. Consequently, EMD black holes are born with an intrinsic, non-trivial scalar field, often referred to as “dilaton hair” and they never bifurcate from a hairless RN seed. Their asymptotic behavior is distinct, approaching a dilaton asymptotic at infinity, and they can violate the RN extremal bound on the charge-to-mass ratio Garfinkle:1990qj ; belkhadria_2025 .
In contrast, the spontaneously scalarized RN branch emerges from a different class of coupling functions satisfying $K(0)=1$, $K^{\prime}(0)=0$, and $K^{\prime\prime}(0)<0$. This specific set of conditions permits the RN metric with $\phi\equiv 0$ as a legitimate and stable background in the absence of a strong electric field. However, in the presence of a sufficiently strong electric charge, a tachyonic instability can arise. The effective mass squared of the scalar field, $\mu_{\text{eff}}^{2}(r)=K^{\prime\prime}(0)\,Q^{2}/r^{4}$, becomes negative within a certain radial window. This instability triggers a second-order phase transition, where the hairless RN black hole spontaneously acquires scalar hair, bifurcating from the original RN branch at a critical charge Qiu_2021 . Asymptotically, these scalarized black holes return to the RN solution, as the scalar hair vanishes at infinity.
Since the EMS theory allows for hairy BH solutions 143 ; 144 ; 145 ; 189 , where BHs can possess scalar hair, investigating their properties in the strong-field regime is of significant interest. Previous studies have shown that dilaton BH spacetimes are influenced by the dilaton charge, which not only serves as a new form of hair but also provides a powerful tool for analyzing optical phenomena in such backgrounds. In this work, we consider a charged BH solution within the EMS theory as an extension of the RN solution. Unlike dilaton BHs, where the scalar field typically contributes through an effective charge, the BH solution in EMS theory is directly governed by its model parameters, which serve as the primary distinguishing feature. A comprehensive exploration of the physical and optical properties of this solution is therefore essential. Investigating the effects of its parameters on spacetime geometry and optical phenomena will contribute to a deeper understanding of its astrophysical implications and provide a means to differentiate it from other BH solutions.
In this article, we analyzed the optical properties of this BH solution through its shadow and the observable appearance of accretion disk emission in the strong-field regime. We further investigated the impact of EMS BH parameters on optical phenomena and constrained the viable ranges of parameters $\alpha$ and $\beta$ using EHT observational data. Our paper is organized as follows. In Sec.II, we will give the BH solution in EMS theory. In Sec.III, we will discuss the influence of EMS parameters on the optical properties of BH and plot the photon orbits of BH. Furthermore, we investigated the observable effects of radiation emitted by the accretion disk and found that the EMS BH exhibits a significantly reduced radiation intensity compared to the Schwarzschild and RN BHs. And in Sec III.E, we also explore the deflection angle in the weak gravitational lensing. Finally, we present our conclusions and provide an outlook in Sec.IV. The paper has been written with the unit system $G=M=c=1$, and metric signature (,+,+,+)(-,+,+,+).
II The black hole solution in Einstein-Maxwell-scalar theory
The systematic investigation of scalar-gravity coupling was introduced by Fisher, who obtained static spherically symmetric solutions to Einstein’s equations with a massless scalar field 102 . This foundational work established the theoretical basis for subsequent developments, particularly in Einstein-Maxwell-scalar (EMS) theory which has attracted significant attention due to its implications and natural emergence in: (i) Kaluza-Klein dimensional reduction 103 , (ii) dilaton couplings in string theory 104 , and (iii) cosmological inflation and late-time acceleration 105 .
The EMS action takes the form
$S=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}\left[R-2(\nabla\phi)^{2}-K(\phi)F^{2}-V(\phi)\right],$ (1)
where $R$ is the Ricci scalar, $\phi$ is the scalar field. $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ represents the electro-magnetic field tensor. The field equations derived from this action yield
$\begin{cases}\nabla_{\mu}[K(\phi)F^{\mu\nu}]=0,\\ \Box\phi=\frac{1}{4}\left[\partial_{\phi}K(\phi)F^{2}+\partial_{\phi}V(\phi)\right],\\ R_{\mu\nu}=2\partial_{\mu}\phi\partial_{\nu}\phi+\frac{1}{2}g_{\mu\nu}V(\phi)+2K(\phi)(F_{\mu\sigma}F^{\sigma}_{~\nu}-\frac{1}{4}g_{\mu\nu}F^{2}).\end{cases}$ (2)
The coupling function $K(\phi)$ and potential $V(\phi)$ determine the theory’s physical content. Some important solutions with different $K(\phi)$ and $V(\phi)$ are given by 3 ; 4 ; 5 ; 6 ; 7 ; 8 . Notable special solutions include
  • RN-de Sitter BH $K(\phi)=1$, $V=2\Lambda$, $\phi=0$,
  • Dilaton BH $K(\phi)=e^{2\phi}$, $V=0$ 1 ; 2 .
Different choices of the coupling function $K(\phi)$ lead to qualitatively distinct physics. Polynomial forms (such as $K(\phi)=1+\alpha\phi^{2}$) preserve the RN branch but allow for spontaneous scalarization above a critical threshold, where the scalar field effectively screens the electromagnetic charge and produces multiple scalarized branches. Exponential couplings (such as $K(\phi)=e^{2\alpha\phi}$) correspond to dilatonic models in which RN solutions no longer exist and charged black holes necessarily carry scalar hair, with modified extremality and thermodynamics. Other bounded or interpolating forms regulate the growth of scalar hair or bridge between EMS-type and dilatonic behavior.
Our investigation focuses on the parameterized coupling function Yu_2021
$K(\phi)=\frac{(\alpha^{2}+1)e^{-2\phi/\alpha}}{(\alpha^{2}+1+\beta)e^{-2\phi(\alpha^{2}+1)/\alpha}+\beta\alpha^{2}},$ (3)
which exhibits two physically significant limits. (i) $\beta\to\infty$: Decoupled electromagnetic sector. (ii) $\beta\to 0$: Einstein-Maxwell-Dilaton (EMD) theory with $K(\phi)=e^{2\alpha\phi}$. (iii) $\beta\to 0$: Einstein-Maxwell-Dilaton (EMD) theory with $K(\phi)=e^{2\alpha\phi}$.
The physical meaning of $\alpha$ is that it measures the strength of the coupling between the dilaton scalar field and the Maxwell field. It first emerges in the EMD limit ($\beta=0$) and is inherited from low-energy string theory, where the dilaton couples universally to gauge fields. Parameter $\beta$ controls the relative contribution of the RN and dilaton branches in the metric. Regarding its origin, it is introduced to combine the RN and dilaton line elements into a single interpolating family, thereby extending the EMD class.
As shown in Ref.9 , these admit BH solution with metric components
$\begin{cases}ds^{2}=-f(r)dt^{2}+f^{-1}(r)dr^{2}+C(r)d\Omega^{2},\\ f(r)=\left(1-\frac{b_{1}}{r}\right)\left(1-\frac{b_{2}}{r}\right)^{\frac{1-\alpha^{2}}{1+\alpha^{2}}}+\frac{\beta Q^{2}}{C(r)},\\ C(r)=r^{2}\left(1-\frac{b_{2}}{r}\right)^{\frac{2\alpha^{2}}{1+\alpha^{2}}},\end{cases}$ (4)
where the parameters $b_{1}$ and $b_{2}$ are determined by
$\begin{cases}b_{1}=\left(1+\sqrt{1-q^{2}(1-\alpha^{2})}\right)M,\\ b_{2}=\frac{1+\alpha^{2}}{1-\alpha^{2}}\left[1-\sqrt{1-q^{2}(1-\alpha^{2})}\right]M,\end{cases}$ (5)
with $q\equiv Q/M$ representing the dimensionless charge-to-mass ratio. Moreover, since obtaining analytic expressions for the inner and outer horizons of the EMS black hole is extremely difficult, we consider two special cases, while the general EMS case is determined numerically. (I) $\alpha=1$: With this definition established, the following relations can be derived Yu_2021
$\begin{cases}r_{+}=M\sqrt{2-q^{2}-\beta q^{2}+\sqrt{\left(2-q^{2}\right)^{2}-4q^{2}\beta}},\\ r_{-}=M\sqrt{2-q^{2}-\beta q^{2}-\sqrt{\left(2-q^{2}\right)^{2}-4q^{2}\beta}}.\end{cases}$ (6)
For analytical convenience, we take $Q>0$ without imposing any loss of generality. The condition ensuring the existence of both the inner and outer horizons can be demonstrated to take the form
$0<q<\sqrt{2},\quad\quad\beta<\frac{\left(2-q^{2}\right)^{2}}{4q^{2}}.$ (7)
Moreover, the condition for the existence of a single horizon can be written as
$0<q<\sqrt{2},\quad\quad\beta=\frac{(2-q^{2})^{2}}{4q^{2}},$ (8)
under which the two horizons coincide, giving rise to an extremal black hole. This demonstrates that an extremal configuration can occur once the coupling constant $\beta$ meets the above criterion, irrespective of how small the charge is. (2) $\beta=0$ (the EMD solution): In this case, the inner and outer horizons are determined by $b_{1}$ and $b_{2}$.
[…]
The horizons are determined by the black hole’s mass $M$, charge $Q$, and the dilaton coupling constant $\alpha$. For two distinct real horizons to exist, the condition $q^{2}\leq 1+\alpha^{2}$ must hold. The equality $q^{2}=1+\alpha^{2}$ corresponds to the extremal case, where the two horizons merge into a single degenerate horizon. In contrast, the condition $q^{2}>1+\alpha^{2}$ eliminates the horizons and exposes a naked singularity. Therefore, when $\alpha\neq 1$ with arbitrary $\beta$, or when $\beta=0$ with arbitrary $\alpha$, a case-by-case analysis is necessary to ensure that naked singularities do not emerge in the black hole solutions. Notably, both situations correspond to special subclasses of the EMS framework. In the following analysis, we confine the parameters to these well-defined ranges to guarantee the existence of black hole solution.
Besides these, the black-hole horizons for all other parameter choices are already plotted in Figure 1. In the figure below, we plot the variation of the metric components of the EMS solution with respect to the coordinate radius $r$.
Figure 1: The above three pictures plot the function $f(r)$ for different values of $\beta$$\alpha$ and $q$.
From the above two pictures, we can see there are two intersection points with $y=0$ which means the existence of the inner and outer horizons of BH. However, in the third case (EMD), we observe that both a single intersection and two intersections can occur. All such parameter choices lie within the admissible ranges discussed above, thereby ensuring the absence of naked singularities. Furthermore, we find that for a fixed $q$, an increase in $\alpha$ causes the function’s minimum value to decrease, whereas an increase in $\beta$ raises the minimum value. This may be attributed to the specific structure of the metric.
III Null geodesics and the Shadow
In this section, we will constrain the parameter of $\alpha$ and $\beta$ by comparing the theoretically defined shadow radius with observational data from the EHT. We begin by considering a static, spherically symmetric spacetime as presented in Ref.177
$ds^{2}=-A(r)dt^{2}+B(r)dr^{2}+C(r)\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right).$ (10)
The functions $A(r)$, $B(r)$, and $C(r)$ satisfy the asymptotic flatness conditions, with $A(r)$ and $B(r)$ $\rightarrow 1$ at spatial infinity while $C(r)\rightarrow r^{2}$. For simplicity, we restrict our analysis to null geodesics in the equatorial plane by fixing the polar angle at $\theta=\pi/2$. Under this configuration, the photon’s Hamiltonian takes the form
$H=\frac{1}{2}g^{ik}p_{i}p_{k}=\frac{1}{2}\left(-\frac{p_{t}^{2}}{A(r)}+\frac{p_{r}^{2}}{B(r)}+\frac{p_{\phi}^{2}}{C(r)}\right).$ (11)
It is worth noting that since we consider EMS BH, the metric functions satisfy $A(r)=f(r)$, $B(r)=f(r)^{-1}$, and $C(r)=r^{2}\left(1-\frac{b_{2}}{r}\right)^{\frac{2\alpha^{2}}{1+\alpha^{2}}}$. Subsequently, the equations of motion for null particles are
[…]
Using Eq.(12), we can derive
[…]
where $\tilde{C}\equiv\frac{d}{dr}\frac{1}{C(r)}=-\frac{2\left(1-\frac{b_{2}}{r}\right)^{-\frac{2\alpha^{2}}{1+\alpha^{2}}}}{r^{3}}-\frac{2\alpha^{2}b_{2}\left(1-\frac{b_{2}}{r}\right)^{-1-\frac{2\alpha^{2}}{1+\alpha^{2}}}}{\left(1+\alpha^{2}\right)r^{4}}$. Setting $H=0$, we have
$\frac{p_{t}^{2}}{f(r)}+\frac{p_{r}^{2}}{f(r)^{-1}}+\frac{p_{\phi}^{2}}{C(r)}=0,$ (14)
and it now follows that
$\frac{dr}{d\phi}=\frac{\dot{r}}{\dot{\phi}}=\frac{C(r)}{f(r)^{-1}}\frac{p_{r}}{p_{\phi}}.$ (15)
Substituting the expressions for $p_{r}$ and $p_{t}=-\omega_{0}$, we finally obtain
$\frac{dr}{d\phi}=\pm\frac{r^{2}}{C(r)^{1/2}f(r)^{-1/2}}\sqrt{\frac{\omega^{2}_{0}}{p^{2}_{\phi}}h(r)^{2}-1},$ (16)
where $h(r)^{2}=\frac{C(r)}{f(r)}$ is defined.
For a circular photon orbit, the conditions $\dot{r}=0$ (vanishing radial velocity) and $\ddot{r}=0$ (zero radial acceleration) necessarily imply $p_{r}=0$, which reduces Eq.(14) to the following form
$0=-\frac{w^{2}_{0}}{f(r)}+\frac{p^{2}_{\phi}}{C(r)}.$ (17)
Furthermore, Eq.(17) can be expressed as
$\dot{p}_{r}=\frac{\omega_{0}^{2}f^{\prime}(r)}{f(r)^{2}}+\tilde{C}p^{2}_{\phi}=0.$ (18)
Using Eq.(17) and (18), we obtain
$p^{2}_{\phi}=C(r)\frac{\omega^{2}_{0}}{f(r)}=-\frac{\omega_{0}^{2}f^{\prime}(r)}{\tilde{C}f(r)^{2}},$ (19)
Combining Eq.(18) and Eq.(19), we give analytic expression for the photon sphere radius of the BH 176 ; 184 (See more details in Appendix A).
$0=\frac{d}{dr}h(r)^{2}.$ (20)
Following Eq.(20), the analytical form of the photon sphere radius $r_{\text{ph}}$ is so complicated. We plot the above equation numerically (See Fig.2), where it shows the values of $\alpha$, $\beta$ and $q$ for the photon sphere to exist. When $q=0.5$, as the parameter $\alpha$ increases, the photon sphere radius essentially increases. Then, for a given value of $\alpha$, we observe an decrease in photon sphere radius as the $\beta$ increases. The second picture is exactly opposite from the first picture. When $\alpha=0$, as the parameter $q$ increases, the photon sphere radius essentially decreases. As $q$ approaches 0, the photon sphere radius asymptotically tends to 3 (Schwarzschild BH). Then, for a given value of $q$, we can observe an increase in photon sphere radius as the $\beta$ decreases. And when $\beta=0$, as the parameter $q$ increases, the photon sphere radius essentially decreases. As $q$ approaches 0, the photon sphere radius asymptotically tends to 3. Then, for a given value of $q$, we can observe an increase in photon sphere radius as the $\alpha$ increases.
Figure 2: The above four pictures plot the photon sphere radius under the different parameters $\alpha$, $\beta$ and $q$.
Subsequently, the definition of the shadow radius must be established. The shadow radius behavior can be investigated by considering light rays traced backward in time from a static observer located at radial coordinate $r_{O}$. As shown in Fig.3, the angle $a$ formed between such light rays and the radial direction is expressed as
$\cot a=\left.\frac{\sqrt{g_{rr}}}{\sqrt{g_{\phi\phi}}}\frac{dr}{d\phi}\right|_{r=r_{\mathrm{O}}}=\left.\frac{\sqrt{B(r)}}{\sqrt{C(r)}}\frac{dr}{d\phi}\right|_{r=r_{\mathrm{O}}}.$ (21)
With the help of Eq.(15), we obtain 176
[…]
where $R$ means the minimum radius that the light ray approaches the center and then goes out (turning point of the trajectory). The shadow boundary is determined by past-oriented light rays asymptotically approaching the unstable circular photon orbits at radius $r_{\mathrm{ph}}$. Consequently, the angular radius $a_{\text{sh}}$ of the shadow is obtained by taking the limit $R\rightarrow r_{\mathrm{ph}}$ in Eq.(22) 184 .
$\sin^{2}a_{\mathrm{sh}}=\frac{h\left(r_{\mathrm{ph}}\right)^{2}}{h\left(r_{\mathrm{O}}\right)^{2}}.$ (23)
Here $h(r)$ is well-defined, $h(r)^{2}=C(r)/A(r)$. Note that the critical value $b_{\text{c}}$ of the impact parameter is connected with $r_{\mathrm{ph}}$ by Ref.176
$b_{\mathrm{c}}=h\left(r_{\mathrm{ph}}\right).$ (24)
Therefore, the Eq.(23) can also be expressed as
$\sin^{2}a_{\mathrm{sh}}=\frac{b_{\mathrm{c}}^{2}}{h\left(r_{\mathrm{O}}\right)^{2}},\quad\text{ or }\quad\sin^{2}a_{\mathrm{sh}}=\frac{b_{\mathrm{c}}^{2}A\left(r_{\mathrm{O}}\right)}{C\left(r_{\mathrm{O}}\right)}.$ (25)
For a general static spherically symmetric asymptotically flat spacetime, Ref.176 demonstrates that when the observer’s radial coordinate $r_{O}$ is sufficiently large, the BH shadow radius approaches $b_{c}$. This radius is determined by the location of the photon sphere and the specific form of the metric function $f(r)$. In our paper, we will obey this consideration.
Figure 3: Example of calculation of light ray emitted from the observer’s position into the past under an angle $a$. The BH horizon and the photon sphere are shown; $r_{\mathrm{ph}}$ is the photon sphere radius. The picture is indicated in Ref.176 .
III.1 Constraints on parameters with the EHT observations of M87* and Sgr A*
We will obtain constraints on the coupling parameters $\alpha$ and $\beta$ using EHT data for M87 and Sgr A, focusing on the non-rotating case. This simplification is justified because
  • M87: General-relativistic magnetohydrodynamic simulations indicate that current EHT observations cannot reliably distinguish between Kerr and dilaton BHs based on shadow morphology alone 111 . The observed shadow size is consistent with $3\sqrt{3}(1\pm 0.17)M$, regardless of spherical or axisymmetric models 112 .
  • Sgr A: The rotation parameter is sufficiently small to negligibly affect the shadow radius 110 .
Based on observational data, the shadow of M87* exhibits an angular diameter of $42\pm 3\ \mathrm{\rm\mu as}$, corresponding to a distance of $16.8$ Mpc from Earth and a central BH mass of $(6.5\pm 0.90)\times 10^{9}\ M_{\odot}$ 113 . In parallel, the EHT’s recent findings for Sgr A* report a shadow size of $48.7\pm 7\ \mathrm{\mu as}$, with the source located at a distance of $8277\pm 33$ pc and hosting a supermassive BH of mass $(4.3\pm 0.013)\times 10^{6}\ M_{\odot}$ 114 ; 115 .
The shadow diameter in mass units can be calculated from the derived BH parameters using the following expression 116
$d_{\mathrm{sh}}=\frac{D\theta}{M}.$ (26)
Hence, the theoretical shadow diameter can be obtained via
$d_{\text{sh }}^{\text{theo }}=2R_{\text{sh}}.$ (27)
Therefore, by using the above expression, we get the diameter of the shadow of M87* and Sgr A*
[…]
Our constraints on the model parameters are shown in Fig.4. The corresponding numerical constraints are further presented in Tab.I (for M87*) and Tab.II(for Sgr A*).
Figure 4: These plots are showing the constraints for different coupling parameters $\alpha$ and $\beta$.
[…]
charge $Q$ upper lower upper lower charge $Q$ upper lower upper lower
0.1 M - - - - 0.1M 99.00 - 45.32 -
0.2 M - - - - 0.2M 24.00 - 10.58 -
0.3 M - - - - 0.3M 10.11 - 4.149 -
0.4 M - - - - 0.4M 5.255 - 1.899 -
0.5 M - - - - 0.5M 3.008 - 0.8592 -
0.6 M - - - 3.441 0.6M 1.790 - 0.2960 -
0.7 M - - - 9.889 0.7M 1.058 - - -
0.8 M - - - 22.59 0.8M 0.5866 - - -
TABLE I. The table lists the values of $\alpha$ and $\beta$ at the $1\sigma$ and $2\sigma$ confidence levels for M87*.
[…]
charge $Q$ upper lower upper lower charge $Q$ upper lower upper lower
0.1 M - - - - 0.1M - - 96.42 -
0.2 M - - - - 0.2M - - 23.36 -
0.3 M - - - - 0.3M - - 9.827 -
0.4 M - - - - 0.4M - - 5.094 -
0.5 M - - - - 0.5M - - 2.905 -
0.6 M - - - - 0.6M - - 1.718 -
0.7 M - - - - 0.7M - - 1.005 -
0.8 M - - - - 0.8M - - 0.5457 -
TABLE II. The table lists the values of $\alpha$ and $\beta$ at the $1\sigma$ and $2\sigma$ confidence levels for Sgr A*.
The variation of the diameter of the shadow image with coupling parameter $\alpha$ and $\beta$ for M87* and for Sgr A* is shown in Fig.4, showing uncertainties at $1\sigma$ and $2\sigma$ levels. The numerical values for the upper or lower bounds in $\beta$ and $\alpha$ is found in Tab.I and Tab.II. However, it turns out that the data for M87* gives a better constraint for the coupling parameter $\beta$. It would mean that there is a certain value for the coupling parameter that gives the observed value of M87* shadow. For the case of Sgr A*, we find that the parameter $\alpha$ remains unconstrained within the $1\sigma$ and $2\sigma$ confidence intervals, whereas the parameter $\beta$ is well constrained within the $1\sigma$ interval.
III.2 Light bending: direct emission, lensing ring and photon ring
To systematically investigate the observational characteristics of EMS BH, we examine their photon ring and lensing ring structures under the assumption of an optically thick accretion disk 177 ; 178 ; 179 . Our analysis begins with studying null geodesics in the EMS BH.
For computational convenience, we employ the coordinate transformation $u\equiv 1/r$, which yields the compact orbital equation
$\left(\frac{du}{d\phi}\right)^{2}=G(u),$ (29)
where the function $G(u)$ takes the explicit form
$\sqrt{G(u)}=\frac{1}{C(1/u)^{1/2}f(1/u)^{-1/2}u^{2}}\sqrt{\frac{\omega^{2}_{0}}{p^{2}_{\phi}}h(1/u)^{2}-1}.$ (30)
Similarly, we restrict our analysis to equatorial motion. The system’s Lagrangian formalism provides two conserved quantities, as derived in Eq.(13). The critical impact parameter $b_{c}$ separates distinct null geodesic behaviors
  • Case 1 ($b>b_{c}$): Photons from asymptotic infinity are gravitationally scattered, reaching a minimum coordinate distance $r_{\text{min}}$ before returning to infinity.
  • Case 2 ($b<b_{c}$): Photons become captured by the BH, demonstrating the existence of a photon capture cross-section.
  • Case 3 ($b=b_{c}$): Photons spiral asymptotically toward the photon sphere at $r_{\text{ph}}$, defining the boundary of BH shadow.
For $b>b_{c}$, the turning point of the trajectory corresponds to the smallest positive real root of $G(u)=0$, denoted as $u_{\rm m}=1/r_{\rm min}$ with $r_{\rm min}$ being the light ray’s radial minimal distance from its trajectory to the BH. Hence, the total change in the azimuthal angle $\phi$ outside the horizon of the trajectory with a given impact parameter $b$ can be expressed as 178 ; 179
$\phi=2\int_{0}^{u_{\rm m}}\frac{du}{\sqrt{G(u)}},\quad b>b_{c}.$ (31)
For $b<b_{c}$, the total change of angle $\phi$ is given by
$\phi=\int_{0}^{u_{0}}\frac{du}{\sqrt{G(u)}},\quad b<b_{c},$ (32)
where $u_{0}\equiv 1/r_{+}$.
Figure 5: The behavior of photon trajectories around the Schwarzschild, RN, and EMS BHs as a function of the impact parameter $b$. In the upper panel, we present the total number of orbits, defined as $n=\phi/2\pi$. The trajectories are categorized based on $n$, where direct emission $n<3/4$ is shown in black, lensed trajectories $3/4<n<5/4$ in yellow, and photon ring trajectories $n>5/4$ in red. The lower panel displays selected photon trajectories in Euclidean polar coordinates $(r,\phi)$. The spacing in the impact parameter is set to $1/10$, $1/100$, and $1/1000$ for the direct, lensed, and photon ring trajectories, respectively. The BH is represented as a solid disk, while the dashed black circle in the ray-tracing diagram marks the photon orbit. For the three cases studied, we set $\alpha=0,\beta=0,q=0$ for the Schwarzschild BH (first column), $\alpha=0,\beta=0,q=0.5$ for the RN BH (second column), and $\alpha=0.5,\beta=0.8,q=0.5$ for the EMS BH (third column).
Following the trajectory classification established in Refs. 120 ; 178 where photon orbits are categorized into direct, lensed, and photon ring components, we quantify the orbital winding through the turn number $n(b)\equiv\phi/2\pi$. This quantity exhibits discrete quantization as demonstrated in Refs. 178 ; 179 , with the functional dependence on impact parameter $b$ given by
$n(b)=\frac{2m-1}{4},\quad m\in\mathbb{Z}^{+}.$ (33)
The critical impact parameters $b_{m}^{\pm}$ demarcate these orbital families, where $b_{m}^{-}$ (satisfying $b_{m}^{-}<b_{c}$) and $b_{m}^{+}$ (satisfying $b_{m}^{+}>b_{c}$) correspond respectively to the minimal and maximal solutions for each $m$-th order orbit, with $b_{c}$ denoting the photon sphere critical parameter. Then the rays can be classified as follows
  • Direct: $\frac{1}{4}<n<\frac{3}{4}\Rightarrow b\in\left(b_{1}^{-},b_{2}^{-}\right)\cup\left(b_{2}^{+},\infty\right)$
  • Lensed: $\frac{3}{4}<n<\frac{5}{4}\Rightarrow b\in\left(b_{2}^{-},b_{3}^{-}\right)\cup\left(b_{3}^{+},b_{2}^{+}\right)$
  • Photon ring: $n>\frac{5}{4}\Rightarrow b\in\left(b_{3}^{-},b_{3}^{+}\right)$
The behaviors of photons in different BHs are plotted in Fig.5. The up panel of Fig.5 depicts the total number of orbits. And the down panel depicts the trajectories of light rays surrounding the BHs, such as ”direct” will intersect the equatorial plane only once.
Parameter $\alpha=0,\beta=0,q=0$ $\alpha=0,\beta=0,q=0.5$ $\alpha=0.5,\beta=0.8,q=0.5$
Direct Emission $b<5.01514$ $b<4.77294$ $b<4.55452$
$n<3/4$ $b>6.16757$ $b>5.97448$ $b>5.81006$
Lensing Ring $5.01514<b<5.18781$ $4.77294<b<4.95793$ $4.55452<b<4.75392$
$3/4<n<5/4$ $5.22794<b<6.16757$ $5.0039<b<5.97448$ $4.80685<b<5.81006$
Photon Ring
$n>5/4$ $5.18781<b<5.22794$ $4.95793<b<5.0039$ $4.75392<b<4.80685$
TABLE III. The region of direct emission, lensing ring, and photon ring for the different parameter $\alpha,\beta,q$.
Table III illustrates the variation of the BH shadow with increasing parameters $\alpha$, $\beta$, and $q$. From the table and corresponding figures, it is observed that the range of the lensing and photon rings expands progressively from the Schwarzschild BH to the RN BH and further to the EMS BH. This indicates that, from left to right, the contribution to the brightness of the lensing and photon rings increases accordingly. Moreover, when the impact parameter approaches the critical value $b\to b_{c}$, the photon orbit exhibits a sharp peak in the $(b,\phi)$ plane. Beyond this regime, as $b$ further increases, the photon trajectories correspond predominantly to direct emission across all cases considered.
III.3 Transfer functions and observed specific intensities
In this subsection, we will derive the specific intensity $I_{\nu}^{\mathrm{em}}$ radiated by the accretion disk, assuming local isotropy of radiation in the rest frame of static observers. The observed intensity can be expressed as 177 ; 178 ; 179
$I_{\nu^{\prime}}^{\mathrm{obs}}=g^{3}I_{\nu}^{\mathrm{em}},$ (34)
where $g=\sqrt{f(r)}$ and $I_{\nu^{\prime}}^{obs}$ are the observed intensity at the frequency $\nu^{\prime}$. By integrating over all frequencies, we obtain the total emitted intensity as $I^{\mathrm{em}}=\int I_{\nu}^{\mathrm{em}}d\nu$. Thus, the corresponding observed frequency can be written as
$I^{\mathrm{obs}}=g^{4}I^{\mathrm{em}}.$ (35)
$I^{\mathrm{em}}$ denotes the integrated specific intensity radiating from the entire accretion disk. Hence, the total received intensity is
$I^{\mathrm{obs}}(b)=\left.\sum_{n}g^{4}I^{\mathrm{em}}(r)\right|_{r=r_{m}(b)},$ (36)
where $r_{m}(b)$ is transfer function (It represents the $m$ intersection points with the equatorial plane). The transfer function fundamentally connects radial coordinates with photon impact parameters, we emphasize that our analysis operates under two key idealizations: (1) no light absorption or reflection by the disk, and (2) no environmental intensity attenuation.
Therefore, the first three transfer functions, which can be obtained by 177 ; 178 ; 179
$\left\{\begin{split}&r_{1}(b)=\frac{1}{u\left(\frac{\pi}{2},b\right)},&b\in\left(b_{1}^{-},\infty\right),\\ &r_{2}(b)=\frac{1}{u\left(\frac{3\pi}{2},b\right)},&b\in\left(b_{2}^{-},b_{2}^{+}\right),\\ &r_{3}(b)=\frac{1}{u\left(\frac{5\pi}{2},b\right)},&b\in\left(b_{3}^{-},b_{3}^{+}\right).\end{split}\right$. (37)
As established in Ref.120 , three distinct observational features emerge:(1) The first transfer function corresponds to the direct disk image;(2) The second forms a significantly demagnified lensing ring from the disk’s far side;(3) The third creates the extremely faint photon ring from the near side. The corresponding graphical representation is presented in the figure below.
Figure 6: The first three transfer functions in BHs for different values of $\alpha,\beta,q$. The above figure, from left to right, respectively Schwarzschild BH, RN BH and EMS BH. Here, the y-axis is the transfer function $r_{m}(b)$ and the x-axis is the impact parameter b.
As shown in Fig.6, the black points $(m=1)$ correspond to the direct emission transfer function and exhibit an approximately constant slope, reflecting the redshifted emission profile from the disk. The yellow points $(m=2)$, associated with the lensing ring, display a rapidly increasing slope with growing impact parameter $b$, indicating that the far-side image is significantly demagnified. The red points $(m=3)$ denote the photon ring, where the slope becomes nearly infinite, resulting in extreme demagnification of the near-side image. Thus, the observed flux is primarily governed by the direct emission, with higher-order contributions $(m>3)$ being negligible.
All three BHs exhibit these characteristics. The critical impact parameters are $b_{c}=5.19615$ for the Schwarzschild BH, $b_{c}=4.96791$ for the RN BH, and $b_{c}=4.76598$ for the EMS BH. This result indicates that the Schwarzschild BH has the largest critical impact parameter, corresponding to the largest shadow radius, while an increase in the model parameters leads to a reduction in the BH’s shadow size.
Additionally, we not only learn the aforementioned characteristics but can also quantify both the morphology (width) and intensity changes of the three ring types. First, we discuss how scalar couplings affect the photon ring. As shown in Table III, we have following calculations. In Schwarzschild BH ($q=0,\alpha=0,\beta=0$), we have $\Delta b_{\text{ph}}=b_{3}^{+}-b_{3}^{-}=5.22794-5.18781=0.04013$. In EMS BH ($\alpha=0.5,\beta=0.8,q=0.5$), we have $\Delta b_{\text{ph}}=b_{3}^{+}-b_{3}^{-}=4.80685-4.75392=0.05293$. The EMS BH shows a net increase of $32\%$ in width. A possible explanation is: the scalar field flattens the effective potential near $r_{\text{ph}}$ (due to $C(r)=r^{2}(1-b_{2}/r)^{\frac{2\alpha^{2}}{1+\alpha^{2}}}$), allowing more near-critical orbits.
For lensing ring adjustment, we now properly treat the two disjoint intervals ($b\in(b_{2}^{-},b_{3}^{-})$ and $b\in(b_{3}^{+},b_{2}^{+})$). In Schwarzschild BH ($q=0,\alpha=0,\beta=0$), we have $\Delta b_{\text{lens1}}=b_{3}^{-}-b_{2}^{-}=5.18781-5.01514=0.17267$, $\Delta b_{\text{lens2}}=b_{2}^{+}-b_{3}^{+}=6.16757-5.22794=0.93963$. Total $\Delta b_{\text{lens}}=1.1123$. In EMS BH ($\alpha=0.5,\beta=0.8,q=0.5$), we have $\Delta b_{\text{lens1}}=b_{3}^{-}-b_{2}^{-}=4.75392-4.55452=0.1994$, $\Delta b_{\text{lens2}}=b_{2}^{+}-b_{3}^{+}=5.81006-4.80685=1.00321$. Total $\Delta b_{\text{lens}}=1.20261$. Compared to a Schwarzschild BH, the EMS BH shows a net increase of $8.1\%$. While the total lensing ring width increases modestly, the inner sub-interval ($\Delta b_{\text{lens1}}$) shows a $15.5\%$ increase due to stronger light bending near the photon sphere.
Finally we consider the direct emission, increasing $\alpha$ leads to a larger $b_{c}$, whereas increasing $\beta$ decreases $b_{c}$. From Table III, the boundaries of the direct emission region exhibit a systematic shift along the impact-parameter axis.
III.4 Observational features of direct emission, photon and lensing rings
As discussed in the previous section, for an observer at infinity, the specific intensity depends solely on the radial coordinate $r$. In our analysis, we consider the following three toy models for the emissivity profile $I^{\mathrm{em}}$177 ; 178 ; 179 ,
$I_{1}^{\mathrm{em}}(r)=\begin{cases}\left(\frac{1}{r-\left(r_{\mathrm{isco}}-1\right)}\right)^{2},&r>r_{\mathrm{isco}}\\ 0,&r\leq r_{\mathrm{isco}}\end{cases}$ (38)
$I_{2}^{\mathrm{em}}(r)=\begin{cases}\left(\frac{1}{r-\left(r_{\mathrm{ph}}-1\right)}\right)^{3},&r>r_{\mathrm{ph}}\\ 0,&r\leq r_{\mathrm{ph}}\end{cases}$ (39)
$I_{3}^{\mathrm{em}}(r)=\begin{cases}\frac{1-\arctan\left(r-\left(r_{\text{isco }}-1\right)\right)}{1-\arctan\left(r_{\mathrm{ph}}\right)},&r>r_{\mathrm{+}}\\ 0,&r\leq r_{\mathrm{+}}\end{cases}$ (40)
where $r_{\text{isco}},r_{\text{ph }}$ and $r_{\mathrm{+}}$ is the innermost stable orbit, photon sphere and event horizon respectively. These three models each exhibit distinct characteristics. The following figures present the results for three distinct types of BHs.
Figure 7: The observational appearance of the thin disk with different emission profiles for $\alpha=0,\beta=0,q=0$ is shown from a face-on perspective. The first row corresponds to the emission profile intensity as described by model 1, the second row represents model 2, and the third row corresponds to model 3, as outlined in Section III.C. In the plots, the emitted and observed intensities, $I^{\mathrm{em}}$ and $I^{\mathrm{obs}}$, are normalized to the maximum value $I^{0}$ of the emitted intensity outside the event horizon.
Figure 8: The observational appearance of the thin disk with different emission profiles for $\alpha=0,\beta=0,q=0.5$ is shown from a face-on perspective. This figure presents the results of the three models corresponding to RN BH.
Figure 9: The observational appearance of the thin disk with different emission profiles for $\alpha=0.5,\beta=0.8,q=0.5$ is shown from a face-on perspective. This figure presents the results of the three models corresponding to EMS BH.
Figures 7, 8, and 9 display the observed emission from thin accretion disks, intensity distributions, and impact parameter mappings for three BH models under different coupling parameters. In each case, the top row corresponds to the first model. The intensity profile (top-left panel) peaks near the critical impact parameter $b_{c}$ and gradually decreases outward. The direct image (top-middle panel) closely follows the intrinsic emission but is affected by gravitational lensing, leading to a sharp boundary and image distortion. Notably, the photon and lensing rings contribute marginally to the overall intensity due to their limited angular extent and lower peak brightness. The shadow image (top-right panel) reveals the lensing ring as a narrow, bright ring just inside the direct image, while the photon ring appears extremely faint and confined to a narrow region, visible only upon magnification. These results highlight the dominant role of direct emission in observed disk images.
In the second model (middle row), the emission intensity peaks near the photon sphere ($r\sim r_{\mathrm{ph}}$) and gradually decreases outward (middle-left panel). The observed intensity profile (middle-center) similarly shows a peak dominated by direct emission from this region, followed by a steady decline. Although the photon and lensing rings contribute modestly, their overlap with direct emission slightly enhances the total signal. Nonetheless, due to strong gravitational redshift and depolarization, their brightness remains faint and confined to a narrow angle. As shown in the shadow image (middle-right panel), direct emission remains the dominant observable feature.
In the third model (bottom row), the emission intensity peaks near the event horizon ($r_{+}$) and gradually declines with increasing radius (bottom-left panel). Due to strong gravitational redshift, much of this radiation is suppressed before reaching a distant observer. Unlike the previous models, the photon and lensing ring emissions overlay the direct radiation; however, their contributions remain minimal because of severe demagnification and narrow angular extent. Consequently, the observed intensity is still primarily determined by direct emission, as clearly seen in the shadow image (bottom-right panel).
Although Fig.7, 8 and 9 show only a few highly idealized cases of thin disk emission near a BH (viewed face on), it illustrates two key points that we believe will hold quite generally for optically thin disk emission: (1) The emission is dominated by the direct emission, with the lensing ring emission providing only a small contribution to the total flux and the photon ring providing a negligible contribution in all cases. (2) For all three models, as we increase the coupling constant $\alpha,\beta$ and $q$, the observed intensity decreases, and it is much less than what we get for the Schwarzschild BH.
III.5 Weak gravitational lensing and magnification of lensed image
Let us now examine the brightness of the image in the weak field limit by using the light’s deflection angle around EMS BH. We will find the deflection angle by using the Gauss-Bonnet theorem (GBT), originally stated as 667 ; 668
$\iint_{D}KdS+\sum_{a=1}^{N}\int_{\partial D_{a}}\kappa_{\mathrm{g}}d\ell+\sum_{a=1}^{N}\theta_{a}=2\pi\chi(D).$ (41)
Consider a freely orientable $2D$ curved surface $S$ with an infinitesimal area element $dS$, which defines a domain $D$ with a Gaussian curvature $K$. The boundary of $D$ is given by $N$ components, $\partial D_{\mathrm{a}}$ (where $\mathrm{a}=1,2,\ldots,N$), along which the geodesic curvature $\kappa_{\mathrm{g}}$ is integrated over the path element $d\ell$ following a positive convention. The jump angles are denoted by $\theta_{\mathrm{a}}$. The Euler characteristic of the domain, $\chi(D)$, is equal to 1 in this case since $D$ is a non-singular region.
It was shown by Ref.666 that in spherically symmetric spacetime admitting asymptotic flatness, the deflection angle can be written as
$\hat{\alpha}=-\iint_{D}K\,dS.$ (42)
Substituting the EMS solution into the above expression and performing the weak-field approximation yields
$\hat{\alpha}\approx\frac{4M}{b}-\frac{3\pi Q^{2}\left(1-\alpha^{2}+\beta\right)}{4b^{2}}.$ (43)
In the context of gravitational lensing, let us consider the following lens equation, expressed in terms of the light angles $\hat{\alpha}$, $\theta$, and $\eta$ Morozova13 ; Bozza2008lens ; Babar21a ; Al_Badawi_2024
$\theta D_{\mathrm{s}}=\eta D_{\mathrm{s}}+\hat{\alpha}D_{\mathrm{ds}}\,$, (44)
where $D_{\mathrm{s}}$ is the distance between the source and observer, $D_{\mathrm{ds}}$ is the distance between the source and lens, and the angular positions of the image and the source are denoted by $\theta$ and $\eta$, respectively. From Eq. (44), the source’s angular position can be expressed as
$\eta=\theta-\frac{D_{\mathrm{ds}}}{D_{\mathrm{s}}}\frac{\hat{\alpha}(\theta)}{D_{\mathrm{d}}}\,$, (45)
where $D_{\mathrm{d}}$ is the distance between the lens and observer. We note the use of the reduced deflection angle $\hat{\alpha}(\theta)=\xi(\theta)/b$, where $b\equiv D_{d}\theta$ is the impact parameter Bozza2008lens . The angular part, known as the Einstein radius $\theta_{E}$, is derived from the spacetime geometry between the source images and corresponds to the radius of the Einstein ring, provided the image forms a ring-like shape. It is given by Morozova13
$\theta_{E}=\sqrt{2R_{s}\frac{D_{ds}}{D_{d}D_{s}}}\,$, (46)
with $R_{s}=D_{d}\theta_{E}$ as the radius of the Einstein ring.
Next, we investigate the magnification of brightness, defined as the ratio of the total brightness of the lensed image ($I_{\mathrm{tot}}$) to the unlensed source brightness ($I_{*}$). The total magnification is given by Babar21a ; Atamurotov2022
$\mu_{\mathrm{tot}}=\frac{I_{\mathrm{tot}}}{I_{*}}=\sum_{k}\bigg|\frac{\theta_{k}}{\beta}\frac{d\theta_{k}}{d\beta}\bigg|,\quad k=1,2,\dots,j\,.$ (47)
For the specific case of two images, the total magnification simplifies to
$\mu_{\mathrm{tot}}=\frac{x^{2}+2}{x\sqrt{x^{2}+4}}\,$, (48)
where $x=\eta/\theta_{E}$ is a dimensionless parameter.
We conducted a numerical analysis of the total magnification for different values of the parameters $\alpha$, $\beta$, and $q$ for EMS BH (Fig  10). Our findings show that for a fixed value of $q$, the total magnification increases with increasing $\alpha$ but decreases with increasing $\beta$. Furthermore, the total magnification decreases as the BH charge increases. We also observed a downward shift in the magnification curves for increasing values of $\beta$ when $\alpha=0$. Conversely, for $\beta=0$, we obtain the opposite trend compared to the $\alpha=0$ case.
Figure 10: Total magnification $\mu_{tot}$ plotted as a function of the different parameters $\alpha$, $\beta$ and $q$.
IV Conclusions
In this paper, we investigated the observational characteristics of thin accretion disk surrounding the EMS BH. We first analyzed the conditions under which the BH possesses inner and outer horizons and derived the theoretical parameter range. Furthermore, we examined the geodesic equation of photons and determined the location of the photon sphere for the EMS BH. Our results indicate that as the parameter $\alpha$ increases, the radius of the photon sphere enlarges, whereas an increase in $\beta$ leads to a decrease in the photon sphere radius. Subsequently, employing the analytical expression for the shadow of spherically symmetric BH, we constrained the model parameters $\alpha$ and $\beta$ using the angular diameter distance from the EHT data within the $1\sigma$ and $2\sigma$ confidence intervals. Our findings reveal that the parameter $\beta$ is well constrained, while the constraints on $\alpha$ remain less stringent.
On the other hand, we analyzed the photon trajectories of the Schwarzschild BH, RN BH, and EMS BH. The differences among these three types of BHs are only reflected in the critical impact parameter $b_{c}$ (the BH shadow radius). We found that the Schwarzschild BH has the largest shadow radius, while the EMS BH has the smallest. All three types of BHs undergo a transition from direct emission to the photon ring, then to the lensing ring, and finally back to direct emission as the impact parameter $b$ increases. A similar effect has been observed using the Okyayvg Mathematical Notebook package 182 in our investigation of BH shadows and rings with three toy models of thin accretion disk. Our analysis reveals that the observed intensity is primarily dominated by direct emission, while the lensing ring contributes only marginally to the total flux, and the photon ring remains negligible in all cases. Among the three types of BHs, the Schwarzschild BH exhibits the highest emission intensity, followed by the RN BH, with the EMS BH having the lowest intensity. In other words, an increase in the model parameters leads to a reduction in the disk’s radiation intensity. Fianlly, when we consider weak gravitational lensing in EMS BH, our results indicate that for fixed $q$, the total magnification grows with $\alpha$ but decreases with $\beta$ and with increasing BH charge. Moreover, $\beta$ shifts the curves downward when $\alpha=0$, while for $\beta=0$ the trend reverses.
Acknowledgements.
This work was supported by the National Key Research and Development Program of China (No. 2022YFA1403700), NSFC (Grants No. 12334002), Guangdong Provincial Quantum Science Strategic Initiative Grand No. SZZX2401001, the Science, Technology and Innovation Commission of Shenzhen Municipality (No. ZDSYS20190902092905285), and Center for Computational Science and Engineering at Southern University of Science and Technology.
Appendix A Derivations for $r_{ph}$
For a general static, spherically symmetric spacetime, the photon sphere radius $r_{\mathrm{ph}}$ can be determined by the extremum condition of the effective potential for null geodesics. Using the metric
$ds^{2}=-f(r)dt^{2}+f^{-1}(r)dr^{2}+C(r)\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right),$ (49)
the condition for the photon sphere reads
$\frac{d}{dr}\left(\frac{C(r)}{f(r)}\right)=0,$ (50)
or equivalently:
$f^{\prime}(r)\,C(r)-f(r)\,C^{\prime}(r)=0.$ (51)
In our EMS black hole model, the metric functions are
$\begin{cases}f(r)=\left(1-\frac{b_{1}}{r}\right)\left(1-\frac{b_{2}}{r}\right)^{\frac{1-\alpha^{2}}{1+\alpha^{2}}}+\frac{\beta Q^{2}}{C(r)},\\ C(r)=r^{2}\left(1-\frac{b_{2}}{r}\right)^{\frac{2\alpha^{2}}{1+\alpha^{2}}},\end{cases}$ (52)
with
$\begin{cases}b_{1}=\left(1+\sqrt{1-q^{2}(1-\alpha^{2})}\right)M,\quad\\ b_{2}=\frac{1+\alpha^{2}}{1-\alpha^{2}}\left[1-\sqrt{1-q^{2}(1-\alpha^{2})}\right]M.\end{cases}$ (53)
Now, for simplicity, we define
$p=\frac{1-\alpha^{2}}{1+\alpha^{2}},\quad s=\frac{2\alpha^{2}}{1+\alpha^{2}},$ (54)
then we have
$\begin{cases}C^{\prime}(r)&=2r\left(1-\frac{b_{2}}{r}\right)^{s}+sb_{2}\left(1-\frac{b_{2}}{r}\right)^{s-1},\\ f^{\prime}(r)&=\frac{b_{1}}{r^{2}}\left(1-\frac{b_{2}}{r}\right)^{p}+\left(1-\frac{b_{1}}{r}\right)p\left(1-\frac{b_{2}}{r}\right)^{p-1}\frac{b_{2}}{r^{2}}-\beta Q^{2}\frac{C^{\prime}(r)}{C(r)^{2}}.\end{cases}$ (55)
Substituting $f(r),f^{\prime}(r),C(r)$ and $C^{\prime}(r)$ into the photon sphere condition
$f^{\prime}(r)\,C(r)-f(r)\,C^{\prime}(r)=0,$ (56)
we obtain the explicit nonlinear equation for $r_{\mathrm{ph}}$ in terms of the model parameters $\alpha,\beta,Q,M$,
$0=$ ${\left[\frac{b_{1}}{r^{2}}\left(1-\frac{b_{2}}{r}\right)^{p}+\left(1-\frac{b_{1}}{r}\right)p\left(1-\frac{b_{2}}{r}\right)^{p-1}\frac{b_{2}}{r^{2}}-\beta Q^{2}\frac{2r\left(1-\frac{b_{2}}{r}\right)^{s}+sb_{2}\left(1-\frac{b_{2}}{r}\right)^{s-1}}{r^{4}\left(1-\frac{b_{2}}{r}\right)^{2s}}\right]r^{2}\left(1-\frac{b_{2}}{r}\right)^{s}}$ (57)
$-\left[\left(1-\frac{b_{1}}{r}\right)\left(1-\frac{b_{2}}{r}\right)^{p}+\frac{\beta Q^{2}}{r^{2}\left(1-\frac{b_{2}}{r}\right)^{s}}\right]\times\left[2r\left(1-\frac{b_{2}}{r}\right)^{s}+sb_{2}\left(1-\frac{b_{2}}{r}\right)^{s-1}\right].$
This equation is highly nonlinear and cannot be solved analytically in closed form for arbitrary parameter values. Therefore, the photon sphere radius can only be expressed implicitly.
However, once specific values of $\alpha,\beta,Q,M$ are chosen, the implicit equation can be solved numerically to yield $r_{\mathrm{ph}}$ with high precision using standard computational tools. This approach allows for direct investigation of the dependence of the photon sphere radius on the physical and coupling parameters of the EMS black hole.
Appendix B Derivations for $\hat{\alpha}$
We follow the Asahi Ishihara’s method 666 to calculate the deflection angle. At first, we have
$dt^{2}=\gamma_{ij}dx^{i}dx^{j},$ (58)
where $i$ and $j$ denote 1,2 and 3 , and $\gamma_{ij}$ is often called the optical metric. Then we can have the Gaussian Curvature
$K=\frac{R_{r\phi r\phi}}{g},\quad g=\det(\gamma_{ij})=\frac{C(r)}{f(r)^{3}},$ (59)
and the Gaussian Curvature then can be written as
$K=-\frac{1}{2}\sqrt{\frac{f^{3}(r)}{C(r)}}\frac{\partial}{\partial r}\left(\sqrt{\frac{f^{3}(r)}{C(r)}}\frac{\partial}{\partial r}\left(\frac{C(r)}{f(r)}\right)\right),$ (60)
thus (in the weak limit $M\ll r$)
$K\approx$ $\frac{1}{r^{6}\left(r^{2}-Mrq^{2}\alpha^{2}\right)^{2}}\left(M(3M-2r)r^{6}+Mr^{2}rq^{2}\left(r^{4}+3M^{3}rq^{2}-2M^{2}r\left(r+rq^{2}\right)\right)\alpha^{2}\right$. (61)
$+Q^{2}(2M-r)r\left(3r^{4}-4Mr^{2}rq^{2}\alpha^{2}+3M^{2}rq^{4}\alpha^{4}\right)\left(-1+\alpha^{2}-\beta\right)$
$\left.+Q^{4}\left(2r^{4}-3Mr^{2}rq^{2}\alpha^{2}+2M^{2}rq^{4}\alpha^{4}\right)\left(1-\alpha^{2}+\beta\right)^{2}\right).$
Next, we will use these equations to find the weak deflection angle using the Gauss-Bonnet theorem (GBT), originally stated as 667 ; 668
$\iint_{D}KdS+\sum_{a=1}^{N}\int_{\partial D_{a}}\kappa_{\mathrm{g}}d\ell+\sum_{a=1}^{N}\theta_{a}=2\pi\chi(D).$ (62)
[…]
It was shown by 666 that in a SSS spacetime admitting asymptotic flatness, the deflection angle can be written as
$\hat{\alpha}=-\iint_{D}K\,dS.$ (63)
Taking the half-plane $\phi\in[0,\pi]$ with area element
$dS\approx r+M\left(3-\frac{a^{2}q^{2}}{2}\right)+\frac{1}{r}\left[-\frac{3}{2}Q^{2}\left(1-a^{2}+\beta\right)+M^{2}\left(\frac{15}{2}-\frac{3a^{2}q^{2}}{2}-\frac{a^{4}q^{4}}{8}\right)\right]dr\,d\phi,$ (64)
integrate over $r\in[r_{0},\infty]$.
The final result is
$\hat{\alpha}\approx\frac{4M}{b}-\frac{3\pi Q^{2}\left(1-\alpha^{2}+\beta\right)}{4b^{2}}.$ (65)